OpenStudy (haleyelizabeth2017):

Write the equation -2x+6y=7 in polar form.

OpenStudy (mathmale):

What are the conversion relationships, relating x to radius r and angle theta, and relating y to radius r and angle theta?

OpenStudy (mathmale):

x=? y=? Write in the proper formulas. Yes, you're right on that x.

OpenStudy (haleyelizabeth2017):

x=r cos theta y=r sin theta

OpenStudy (haleyelizabeth2017):

Which I then got -2 r cos theta+6 r sin theta=7, I just wasn't sure how to continue on

myininaya (myininaya):

you could solve for r

myininaya (myininaya):

but you have write it already in polar form

myininaya (myininaya):

but if you wanted to solve for r you could factor out the r from the two terms on the left hand side and then divide by what is being multiplied by r on both sides

OpenStudy (haleyelizabeth2017):

Would r(cos theta+sin theta)=-7/12 be correct?

myininaya (myininaya):

what happen to -2 next to cos(theta) and the 6 next to the sin(theta) also where did -1/12 come from

OpenStudy (haleyelizabeth2017):

Wait...

OpenStudy (haleyelizabeth2017):

Sorry, I was confused by what you were having me do, but I think I understand now..

OpenStudy (haleyelizabeth2017):

\[r=\frac{7}{(-2cos\theta+6sin\theta)}\]

myininaya (myininaya):

looks great

OpenStudy (haleyelizabeth2017):

What would I do next?

myininaya (myininaya):

nothing really you can do other things you can do are just up to the viewer like some math people don't like the term with - sign first they might prefer to write this...since it cost less amount of symbols: \[r=\frac{7}{6 \sin(\theta)-2 \cos(\theta)}\]

OpenStudy (haleyelizabeth2017):

Well, my question is having me write it like http://imgur.com/Xd97J2h

myininaya (myininaya):

\[\sqrt{6^2+(-2)^2}=\sqrt{6^2+4^2}=\sqrt{36+16}=\sqrt{52}=\sqrt{4 \cdot 13}=2 \sqrt{13} \\ \text{ so anyways} \\ 2 \sqrt{13}(\frac{6}{2 \sqrt{13}} \sin(\theta)-\frac{2 }{2 \sqrt{13}} \cos(\theta)) =... \text{ one sec }...\]  so we have \[2 \sqrt{13}(\frac{6}{2 \sqrt{13}} \sin(\theta)+\frac{-2}{2 \sqrt{13}} \cos(\theta)) \\ =2 \sqrt{13}(\cos(u) \sin(\theta)+\sin(u)\cos(\theta)) \\ = 2 \sqrt{13}( \text{ hint: use a sum identity here })\]